OpenStudy (anonymous):

a recipe for a pie calls for 1/3 cup of blackberries and 2/6 cup of raspberries. What is the total amount of berries needed for the recipe?

OpenStudy (anonymous):

\[\frac{1}{3}+\frac{2}{6}=\frac{1}{3}+\frac{1}{3}=...\]

OpenStudy (anonymous):

^ okay, thats the step i did XD

OpenStudy (anonymous):

there is no recipe that calls for \(\frac{2}{6}\) of anything, since \(\frac{2}{3}=\frac{1}{3}\) they really stretch to make these word problems, don't they?

OpenStudy (anonymous):

it's apex work -.- .

OpenStudy (anonymous):

so there isn't a answer

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

what is \(\frac{1}{3}+\frac{1}{3}\) ?

OpenStudy (anonymous):

2/6 ?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

No... When the denominators are the same, you only add the numerator

OpenStudy (anonymous):

add the numerators, don't touch the denominators

OpenStudy (anonymous):

ohhhhhhhhhhhhhhh hahahah omg . okay so its 2/3

OpenStudy (anonymous):

one apple plus one apple is two apples one third plus one third is two thirds

OpenStudy (anonymous):

yes, that is it

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

so the answer is 2/3 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have another one and many more

OpenStudy (anonymous):

time to learn to add and subtract and reduce fractions then, huh?

OpenStudy (anonymous):

i am sure between myself and @whalexnuker we can help you out

OpenStudy (anonymous):

Totally :P

OpenStudy (anonymous):

You have 3/6 cup of brown sugar in your cupboard. The recipe for a dessert calls for 1/5 cup of brown sugar. How much brown sugar would you have left if you made the dessert?

OpenStudy (anonymous):

can you reduce \(\frac{3}{6}\) first?

OpenStudy (anonymous):

it's okay to use a calculator ;)

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

good

OpenStudy (anonymous):

now \[\frac{1}{2}-\frac{1}{5}=\frac{5\times 1-2\times 1}{2\times 5}\] will get the answer

OpenStudy (anonymous):

whalexnuker i dont know how to reply to your message :o .

OpenStudy (anonymous):

Just hit reply next to it XD its that little blue arrow.

OpenStudy (anonymous):

3/10 >.<

OpenStudy (anonymous):

you can start a new question

OpenStudy (anonymous):

okay , a recipe calls for 2/5 cup of chicken broth and 1/6 cup of vegetable broth. What is the total amount of broth needed for the recipe?

OpenStudy (anonymous):

once again add \[\frac{2}{5}+\frac{1}{6}\] via \[\frac{2\times 6+5\times 1}{5\times 6}\]

OpenStudy (anonymous):

let me know what you get this is how you add fractions

OpenStudy (anonymous):

17/30 ? i dont think that is right though .

OpenStudy (anonymous):

i am pretty sure it is

OpenStudy (anonymous):

in fact, i am positive it is

OpenStudy (anonymous):

okaay sweeeeet (: . thank you (:

OpenStudy (anonymous):

not sure why you doubt it that is how you add fractions you probably need to remember this for future reference

OpenStudy (anonymous):

It's all a matter of the wording of the word problem. A calculator is used at it's best :P